The flower garden has the shape of a right triangle. 5 ft of a perennial border forms the hypotenuse of the​ triangle, and one leg is 1 ft longer than other leg. Find the lengths of the legs.
Answer by jorel555(1290)

Let n be the length of the shortest leg of the triangle. Then the other leg is n+1; and we know the hypotenuse is 5, so:
(n)²+(n+1)²=5²
2n²+2n-24=0
n²+n-12=0
(n+4)(n-3)=0
n=3 or -4
Throwing out the negative result, we get the length of the shortest leg to be 3 ft., and the longer leg to be 4 ft. ☺☺☺☺